# HypeRL: Parameter-Informed Reinforcement Learning for Parametric PDEs

### v1

@@ 4 Numerical Experiments
Figure 6: Training and evaluation results. The solid line represents the mean and the shaded area the standard deviation over 5 different random seeds.
Similarly to the KS case, the results, especially in early stages of the training (see Figure 6 (d)), show the superior performance of the parameter-informed HypeRL-TD3 agents. Again, the agent without access to the parameter $\mu$ achieves the worst performance. The analysis of the confidence interval is consistent with the previous results and show very clearly that, especially in early stages of the training, the HypeRL agents can learn good policies with limited training time.
Figure 7: $95\%$ confidence intervals of the training and evaluation reward for the different phases of training and evaluation. This metric, suggested in [61], allows to assess the reliability of the results accounting for the stochastic nature of the RL experiments.
Eventually, we show controlled solutions for $t=t_{f}$ and $\mu=0.077$ (unseen value inside the training range) in Figure 8 and $\mu=0.11$ (unseen value outside the training range) in Figure 9 and we report the state cost $c_{1}$, and action cost $\alpha c_{2}$.
Figure 8: Controlled solutions $\bm{y}(t=t_{f})$ for $\mu=0.077$.
Figure 9: Controlled solutions $\bm{y}(t=t_{f})$ for $\mu=0.11$.
Analogously to the KS case, the TD3 agent without access to the PDE parameter $\mu$ is the worst-performing method in interpolation and extrapolation regimes. In the two examples shown, the HypeRL-TD3 with only access to the parameter is the best performing agent, closely followed by HypeRL-TD3, and TD3 with knowledge of $\mu$. It is worth highlighting that the differences in performance among the agents is less pronounced than in the KS case. This happens for two main reasons: (i) the control only affects the system through one boundary, leaving the agent less capabilities to freely steer the PDE towards the desired state, and (ii) the reference state and control used in the reward function in Equation (25), provided by PDEControlGym, were generated for the case of $\mu=0.1$, making the control problem extremely complex for different values of the parameters.
5 Conclusion
@@
In this paper, we proposed a novel approach for optimal control of parametric PDEs using RL and hypernetworks that we named HypeRL. In particular, due to the high computational cost of solving PDEs, we focused on developing a framework for enhancing the performance of RL algorithms in terms of sample efficiency and generalization to unseen instances of the PDE parameters when trained on a limited number of instances of such parameters. We tested the capabilities of HypeRL on two challenging control problems of chaotic and parametric PDEs, namely a 1D Kuramoto Sivashinsky and a 2D Navier Stokes. We showed that knowledge of the PDE parameters and how this information is encoded, i.e., via hypernetorks, is an essential ingredient for sample-efficient learning of control policies that can generalize effectively.
@@ Acknowledgments
AM and SF are members of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM) and acknowledge the project “Dipartimento di Eccellenza” 2023-2027, funded by MUR. SF acknowledges the Istituto Nazionale di Alta Matematica (INdAM) for the financial support received through the “Concorso a n. 45 mensilità di borse di studio per l’estero per l’a.a. 2023-2024”. AM acknowledges the PRIN 2022 Project “Numerical approximation of uncertainty quantification problems for PDEs by multi-fidelity methods (UQ-FLY)” (No. 202222PACR), funded by the European Union - NextGenerationEU, and the Project “Reduced Order Modeling and Deep Learning for the real- time approximation of PDEs (DREAM)” (Starting Grant No. FIS00003154), funded by the Italian Science Fund (FIS) - Ministero dell’Università e della Ricerca. AM and SF acknowledge the project FAIR (Future Artificial Intelligence Research), funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence).
Appendix A Neural Network Architectures
A.1 HypeRL Architecture
The main policy and value function networks are composed of one input layer of dimension equal to the state dimension $|\bm{y}_{k}|$, and state and action dimension $|\bm{y}_{k}|+|\bm{u}_{k}|$, respectively, one hidden layer with $256$ neurons, and an output layer of dimension equal to the control action dimension $|\bm{u}_{k}|$, and dimension 1, respectively. These weights and biases are learned by two hypernetworks with inputs the PDE parameter $\bm{\mu}$ or state and PDE parameter $\bm{y}_{k}$ and $\bm{\mu}$. The hypernetworks use the same architecture and weight initialization proposed in [38].
A.2 TD3 Architecture
The TD3 algorithms use the default and widely-used architectures introduced in [58]. The architectures are analogous to the main policy and value function described in Section A.1 but with an additional hidden layer of dimension $256$. The input to these networks are state and PDE parameter, and state, action, and PDE parameter, respectively. The TD3 without parameter has the very same architecture but without access to the PDE parameter.
A.3 Convolutional Encoder 2D Navier-Stokes Equations
Due to the high dimensionality of the state ($21\times 21\times 2=882$ dimensions) in the NS test case, we utilize a convolutional encoder to reduce the dimensionality of the state to $20$ dimensions before feeding it to the RL networks. The convolutional encoder is composed of 4 2D convolutional layers with $32$ output channels, kernel size of $3\times 3$, and stride length equal to $1$ (with the exception of the first layer that uses stride length equal to $2$). After the second and forth convolutional layer, we use a batch-normalization layer. The features are then flattened and fed to two fully-connected layers to reduce the state dimension to $20$. A similar architecture was employed in [33].
Appendix B Additional Results
B.1 1D Kuramoto-Sivashinsky PDE
@@ Appendix B Additional Results
In Figure 10, we show the state and action costs over training and evaluation for the KS PDE.
@@
Figure 10: Training and evaluation results. The solid line represents the mean and the shaded area the standard deviation over 5 different random seeds.
B.2 2D Navier-Stokes Equations
@@
In Figure 11, we show the state and action costs over training and evaluation for the NS equations.
@@
Figure 11: Training and evaluation results. The solid line represents the mean and the shaded area the standard deviation over 5 different random seeds.

### v2

@@ 4.1.2 Stabilization of the State of a Parametric KS Equation to an Arbitrary Reference
Figure 6: Stabilization of the state of a parametric KS equation to an arbitrary reference – training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Table 2, we show the training and evaluation rewards collected by the different agents over 5 different random seeds, where we highlight in bold the highest scores of the training and evaluation. In the calculation of the scores, we remove for every agent the $100$ warmup episodes.
Table 2: Mean and standard deviation of the cumulative reward over training (left) and evaluation (right) collected by the different algorithms. The results report the average performance over 5 different random seeds.
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-223.70\pm 74.20}$
TD3 $396.56\pm 133.33$
TD3 (Huber) $329.48\pm 94.98$
TD3 (2 layers) $450.86\pm 188.58$
TD3 (no $\bm{\mu}$) $4598.64\pm 2294.87$
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-210.50\pm 18.04}$
TD3 $395.40\pm 92.77$
TD3 (Huber) $326.36\pm 30.47$
TD3 (2 layers) $446.20\pm 138.94$
TD3 (no $\bm{\mu}$) $4222.67\pm 485.95$
Again, the results show that the hypernetwork-based algorithm outperforms the TD3 agents, showing that encoding parametric information via a hypernetwork is crucial for sample efficiency and generalization of the RL algorithms.
In Figure 7, we show an example of controlled solution for $\bm{y}_{\text{ref}}=-\bm{1}$ and $\nu=0.25$, where the first row depicts the evolution of the state, the second one the control, and the third one the state-tracking error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
Figure 7: Controlled solutions of the parametric KS equation when using the different agents. The first row shows the state, the second the control, and the third one the error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
Additional results are reported in Appendix B.2.
4.2 Navigation of a Particle in a Gyre Flow
As a second test case, we consider the navigation problem of a mobile particle in a double gyre flow field [21, 30]. The flow field is defined as:
$\phi(x,y,t)=A\sin{\big(\pi f(x,t)\big)}\sin{\big(\pi y\big)}\,$,
where $x,y$ are the spatial coordinates, $t$ the time, $\pi$ (with a slightly abuse of notation) represents pi, and
$f(x,t)=\big[\epsilon\sin(\omega t)\big]x^{2}+\big[1-2\epsilon\sin(\omega t)\big]x\,$,
Hence, an ODE system governing the particle dynamics in this gyro flow field is given by:
[…]
in which $u_{x},u_{y}$ are the control inputs. The flow is restrained within a domain $(x,y)^{T}\in\mathcal{D}=[0,2]\times[0,1]$ and $A,\epsilon,\omega$ are model parameters controlling the flow pattern. In particular, $A$ affects the velocity magnitude, $\epsilon$ the magnitude of the oscillation in the $x$-direction, and $\omega$ the angular oscillation frequency. Similarly to [30], we solve Equation (16) for $T=80$s and $\Delta t=0.1$s. However, differently from [21, 30] where the agent is trained to reach a single target, we consider two more challenging control problems:
  1. 1.
    navigation of a particle from a random initial position to a random target in a gyre flow with fixed parameters, namely $A=0.1,\omega=2\pi/10$, and $\epsilon=0.25$,
  2. 2.
    navigation of a particle from a random initial position to a random target in a parametric gyre flow with $A\in\mathcal{U}(0.1,0.4),\omega\in\mathcal{U}(2\pi/10,2\pi/2)$, and $\epsilon=0.25$.
To force the controller to exploit the dynamics of the gyre flow to navigate as the largest flow velocity is $\approx A\pi$, we limit $u_{x},u_{y}\in[-0.2,0.2]$. In both cases, the agent’s state $\bm{z}_{k}=[x_{k},y_{k},k,\bm{\mu}]$ in which the collection $x_{k},y_{k},k]$ serves as $\bm{y}_{k}$ from Section 3, with $\bm{\mu}=[x_{\text{target}},y_{\text{target}},A,\omega]$, and $x_{\text{target}},y_{\text{target}}$ are the coordinates of the target location. In Figure 8, we show examples of particle trajectories for different values of the amplitude parameter.
[…]
Our goal is to steer a particle in a gyre flow, defined as in Equation (16), from a randomly sampled initial condition $\in\mathcal{U}(0.1,1.9)\times\mathcal{U}(0.1,0.9)$ to a randomly sampled target position $(x_{\text{target}},y_{\text{target}})\in\mathcal{U}(0.1,1.9)\times\mathcal{U}(0.1,0.9)$ with minimal control effort. Therefore, we utilize the following reward function:
[…]
where $\alpha=0.1$ is a scalar positive coefficient balancing the contribution of the action cost.
To assess the effectiveness of our control strategy in exploiting the flow, we introduce the Lyapunov exponents [64, 51, 48]. The Lyapunov exponents (LE) play a crucial role in the investigation of the behaviour of chaotic dynamical systems. LEs measures the rate of convergence or divergence of a trajectory when the initial condition is perturbed. The finite-time Lyapunov exponent (FTLE) generalizes the LE to finite times, i.e., it measures the sensitivity of a trajectory to its initial condition over a finite-time horizon. Through the FTLE of the flow field, we are capable of identifying the areas of the domain where passive particles get trapped in and their ridges, separating the different stable regions. We leave the numerical computations of the FTLE to Appendix A. In general, regions with high FTLE requires higher control effort than regions with lower FTLE, making the latter more favorable for energy-efficient control. Examples of FTLE fields for different values of the parameters $A$ and $\omega$ are shown in Figure 9. The FTLE fields are computed for $T=10$s, $dt=0.1$s, $N_{x}=300$, and $N_{y}=150$. It is worth highlighting that small variations of the gyre flow parameters produce large variations in the FTLE fields, making the learning of optimal control policies extremely challenging.
Figure 9: Finite-time Lyapunov exponent field calculated for $T=10$s and different values of the amplitude AA and frequency $\omega$ parameters - $A=0.1,\omega=2\pi/10$ (top left), $A=0.2,\omega=2\pi/10$ (bottom left), $A=0.3,\omega=2\pi/10$ (top right), and $A=0.1,\omega=2\pi/6$ (bottom right).
4.2.1 Navigation of a Particle to Arbitrary Targets in a Gyre Flow
In this first example, the goal is to control a particle in a gyre flow to a randomly sampled target location inside the domain. In Figure 10, we show the mean and the standard deviation of the cumulative reward during training and evaluation, respectively. It is possible to notice the superior performance of the hypernetwork-based agent in early and later stages of the training and during evaluation.
Figure 10: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Table 3 we report the cumulative rewards over training and evaluation of the different agents, averages over 5 different random seed. The HypeRL-TD3 agent outperforms all the other agent by achieving the lowest average reward upon training and evaluation.
Table 3: Mean and standard deviation of the cumulative reward over training (left) and evaluation (right) collected by the different algorithms. The results report the average performance over 5 different random seeds.
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-40.08\pm 18.45}$
TD3 $102.5\pm 60.46$
TD3 (Huber) $57.77\pm 26.708$
TD3 (2 layers) $131.81\pm 71.74$
TD3 (no $\bm{\mu}$) $344.34\pm 131.50$
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-39.16\pm 6.00}$
TD3 $114.04\pm 41.10$
TD3 (Huber) $57.55\pm 10.82$
TD3 (2 layers) $131.48\pm 25.24$
TD3 (no $\bm{\mu}$) $346.58\pm 25.93$
Eventually, in Figure 11 and 12, we show the trajectory of the particle when controlled by the different agents where the background represents the flow field at the last timestep of the trajectory and the FTLE computed over an interval of $T=10$s with a $dt=0.1$, respectively. While the different variants of TD3 get influenced by the ridges of the FTLE (see Figure 12) and achieve trajectories that oscillate around the target that require more time and control effort, HypeRL-TD3 smoothly and efficiently reaches the target by exploiting the stable regions of the gyre flow field.
Figure 11: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 12: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents overlaid with the FTLE. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Additional results, including the performance of the TD3 agent without access to the knowledge of $\bm{\mu}$, are shown in Appendix C.1.
4.2.2 Navigation of a Particle to Arbitrary Targets in a Parametric Gyre Flow
In the last test case, the goal is to control a particle in a parametric gyre flow. We use the same settings as the previous case with the addition of randomly sampling the parameter $A\in\mathcal{U}(0.1,0.4)$ and the parameter $\omega\in\mathcal{U}(2\pi/10,2\pi/2)$ at the beginning of each training and evaluation episode.
In Figure 13, we show the mean and the standard deviation of the cumulative reward during training and evaluation, respectively. It is possible to notice the superior performance of the hypernetwork-based agent in early and later stages of the training and during evaluation.
Figure 13: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
Table 4: Mean and standard deviation of the cumulative reward over training (left) and evaluation (right) collected by the different algorithms. The results report the average performance over 5 different random seeds.
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-112.38\pm 70.55}$
TD3 $398.12\pm 332.63$
TD3 (Huber) $147.08\pm 109.23$
TD3 (2 layers) $801.62\pm 576.11$
TD3 (no $\bm{\mu}$) $457.47\pm 266.10$
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-110.06\pm 16.05}$
TD3 $430.34\pm 153.73$
TD3 (Huber) $152.05\pm 43.72$
TD3 (2 layers) $942.89\pm 245.60$
TD3 (no $\bm{\mu}$) $441.75\pm 84.12$
Eventually, in Figure 14 and 15, we show the trajectory of the particle when controlled by the different agents where the background represents the flow field at the last timestep of the trajectory and the FTLE computed over an interval of $T=10$s with a $dt=0.1$, respectively. While the different variants of TD3 get influenced by the ridges of the FTLE (see Figure 15) and achieve trajectories that oscillate around the target that require more time and control effort, HypeRL-TD3 smoothly and efficiently reaches the target by exploiting the stable regions of the gyre flow field.
Figure 14: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 15: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents overlaid with the FTLE. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
5 Discussion and Conclusion
@@
In this paper, we proposed a novel approach for optimal control of parametric dynamical systems using RL and hypernetworks that we named HypeRL. In particular, we focused on developing a framework for enhancing the performance of RL algorithms in terms of sample efficiency and generalization to new instances of the dynamical system parameters. We tested the capabilities of HypeRL on two challenging control problems of chaotic and parametric dynamical, namely a 1D Kuramoto Sivashinsky and a navigation problem in a 2D gyre flow. We showed that knowledge of the dynamical systems parameters and how this information is encoded, i.e., via hypernetworks, is an essential ingredient for sample-efficiency and for learning of control policies that can generalize effectively.
In this work, we assumed perfect knowledge of the dynamical system and task-dependent parameters. While the knowledge of the physical parameters might seem a strong assumption, the advances of machine learned has opened the possibility of estimating the physical parameters from raw data, e.g., in [55]. As future work, we plan to incorporate a parameter-estimation step in our control pipeline and develop a two-stage procedure to avoid the need for exact knowledge of the physical parameters.
In our numerical approaches, we showed that the parametrization of policy and value function matters for improving the performance of the agent. Starting from the TD3 agent with three fully connected layers, we tested a two-layer architecture to match the HypeRL-TD3. However, removing one layer and reducing the learnable parameters of the TD3 agent drastically reduces the cumulative reward achieved by the agent. In addition, we replace the mean-squared error loss with the Huber loss to update the value function to match the HypeRL-TD3 update, but while it was beneficial for the improving mean reward and reducing the standard deviation across the different seed compared to TD3, it is still not enough to achieve comparable performance to HypeRL-TD3. These results strengthen our claim that hypernetwork-based parametrizations better encode state, task-dependent, and physical parameters into the learnable parameters of the main policy and value function networks. Although the total learnable parameters are higher due to the presence of hypernetworks, the optimization of these learnable parameters becomes simpler, requires less data, and allows the learning of better controllers.
@@ Acknowledgments
NB and AM acknowledge the Project “Reduced Order Modeling and Deep Learning for the real-time approximation of PDEs (DREAM)” (Starting Grant No. FIS00003154), funded by the Italian Science Fund (FIS) - Ministero dell’Università e della Ricerca. AM also acknowledges the project “Dipartimento di Eccellenza” 2023-2027 funded by MUR and the project FAIR (Future Artificial Intelligence Research), funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence). NB and AM are members of the National Group of Scientific Computing (GNCS).
Appendix A Numerical Computation of the Finite Time Lyapunov Exponent
To compute the FTLE field for the dynamical systems $\bm{F}(\bm{y}(t))$ we follow the same procedure as in [30]. We initialize the grid of passive particles at time $t_{0}=0.0$ and integrate them using $\bm{F}(\bm{y}(t))$ for a fixed amount of time $T=10.0$. This results in a flow map:
$\bm{\Phi}_{t_{0}}^{t_{0}+T}:\bm{y}(t_{0})\mapsto\bm{y}(t_{0})+\int_{t_{0}}^{t_{0}+T}\bm{F}(\bm{y}(\tau))d\tau.$ (18)
The flow map operator $\bm{\Phi}_{t_{0}}^{t_{0}+T}$ maps an initial state at $t=t_{0}$ to a final state at $t=t_{0}+T$. After computing the flow map, its Jacobian can be computed using finite differences as:
$(\bm{D}\bm{\Phi}_{t_{0}}^{t_{0}+T})_{i,j}=\Bigg[\begin{matrix}d_{0,0}&d_{0,1}\\ d_{1,0}&d_{1,1}\\ \end{matrix}\Bigg]$ (19)
where $i,j$ correspond to the indices of the particles in the grid and
[…]
where $\bm{y}=[x,y]$. Using the Jacobian of the flow map, we can compute the Cauchy-Green deformation tensor
$\bm{\Delta}_{i,j}=(\bm{D}\bm{\Phi}_{t_{0}}^{t_{0}+T})_{i,j}^{\top}(\bm{D}\bm{\Phi}_{t_{0}}^{t_{0}+T})_{i,j}\,$, (21)
where $(\cdot)^{\top}$ represents the matrix transpose. Eventually, we can compute compute the FTLE field from $\bm{\Delta}_{i,j}$ for each particle as:
$\sigma_{i,j}=\frac{1}{T}\ln\sqrt{(\lambda_{\max})_{i,j}}\,$, (22)
where $\lambda_{\max}$ indicates the largest eigenvalue.
Appendix B Additional Results: Stabilization of a Kuramoto-Sivashinsky Equation
B.1 Stabilization of the State of a KS Equation to an Arbitrary Reference
@@
In Figure 16, we show the state and action costs over training and evaluation for the KS equation.
@@
Figure 16: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Figure 17 and 18, we show examples of controlled solution when tracking $\bm{y}_{\text{ref}}=-\bm{1}$ and $\bm{y}_{\text{ref}}=\bm{0}$.
Figure 17: Controlled solutions of the KS equation when using the different agents. The first row shows the state, the second the control, and the third one the error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
Figure 18: Controlled solutions of the KS equation when using the different agents for $\bm{y}_{\text{ref}}=\bm{0}$.
B.2 Stabilization of the State of a Parametric KS Equation to an Arbitrary Reference
@@
In Figure 19, we show the state and action costs over training and evaluation for the KS equation.
@@
Figure 19: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Figure 20 and 21, we show examples of controlled solution when tracking $\bm{y}_{\text{ref}}=-\bm{1}$ and $\nu=0.25$ and $\bm{y}_{\text{ref}}=\bm{2}$ for $\nu=-0.2$.
Figure 20: Controlled solutions of the parametric KS equation when using the different agents. The first row shows the state, the second the control, and the third one the error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
Figure 21: Controlled solutions of the KS equation when using the different agents. The first row shows the state, the second the control, and the third one the error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
Appendix C Additional Results: Particle Navigation in a Gyre Flow
C.1 Navigation of a Particle to Arbitrary Targets in a Gyre Flow
In Figure 22, we show the state and action costs over training and evaluation for the gyro flow test case.
Figure 22: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Figures 23 and 25 we show examples of controlled trajectories overlaid with the gyro flow field, while in Figures 24 and 26 we overlaid the trajectories with the FTLE.
Figure 23: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 24: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents overlaid with the FTLE. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 25: Trajectories of a controlled particle in a gyre flow when using the different agents.
Figure 26: Trajectories of a controlled particle in a gyre flow when using the different agents.
C.2 Navigation of a Particle to Arbitrary Targets in a Parametric Gyre Flow
In Figure 27, we show the state and action costs over training and evaluation for the gyro flow test case.
Figure 27: Training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
In Figures 28 and 30 we show examples of controlled trajectories overlaid with the gyro flow field, while in Figures 29 and 31 we overlaid the trajectories with the FTLE.
Figure 28: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 29: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents overlaid with the FTLE. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 30: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.
Figure 31: Trajectories (magenta dash line) of a controlled particle (yellow hexagon) in a gyre flow when using the different agents overlaid with the FTLE. The magenta circle indicates the starting location and the magenta cross the target one and the blue arrows represent the control inputs.